OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. 2x - y = 8 y = x - 3

OpenStudy (anonymous):

replace y by \(x-3\) in the first equation and solve \[2x-(x-3)=8\] for x

Parth (parthkohli):

It's a very easy method. We have to replace the value. By the 2nd line, we can infer that 2x - (x - 3) = 8 \[2x - x + 3 = 8\] \[x + 3 = 8\] \[x = 5\] We know that y is x-3, so y = 2